Circular No. 6090

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

PERIODIC COMET MACHHOLZ 2 (1994o) P. Pravec, Ondřejov Observatory, reports that his CCD observations

with the 0.65-m f/3.6 reflector on Oct. 5.14 UT showed component D to consist of two condensations of similar brightness and separated by 7′′ in p.a. 112^{o}–292^{o}. Component C was marginally visible at separation 10′′ in p.a. 185^{o} but almost hidden by the coma of component D. Component E was also visible at separation ∼ 80′′ in p.a. 9^{o} from component D. There was also a trail of material along a line in p.a. 10^{o}–190^{o} near component D; this spanned from 2′.0 at the southern end to 5′ at the northern and was measured on a composite image of 240-s total integration (although it was also visible on individual 60-s images). The duplicity of component D is also suggested on Sept. 2.1 and 4.1, when the single condensation was elongated to about the same extent and in the same direction as the resolved condensations are now. The elongation is less prominent, but still present, in images from Aug. 30.1. On Sept. 23.1 component D was significantly more active and better condensed than on other occasions.

Total visual magnitude estimates: component A, Oct. 2.49 UT, 9.1 (A. Hale, near Chaparral, NM, 0.20-m reflector); 3.49, 9.2 (C. S. Morris, southern California, 0.26-m reflector); 4.15, 9.2 (J. D. Shanklin, Cambridge, England, 0.33-m reflector); 5.45, 9.8 (G. Kronk, Troy, IL, 0.33-m reflector); component B, Sept. 24.52 UT, 12.2 (Morris); Oct. 3.50, 12.5: (Morris); component D, Oct. 2.49, 9.5 (Hale); 3.49, 9.8 (Morris); 4.15, 10.4 (Shanklin); 5.44, 11.4 (Kronk); component E, Sept. 11.50 UT, 12.2 (Morris); 24.51, 12.8 (Morris); Oct. 3.50, 12.8 (Morris).

SUPERNOVA 1994aa IN NGC 1320 A. Riess and R. Kirshner, Harvard-Smithsonian Center for Astro-

physics, communicate: “A spectrogram obtained with the Multiple Mirror Telescope (+ Blue Channel spectrograph, range 400–800 nm) on Oct. 5 shows that SN1994aa is a type Ia supernova ∼ 1 month past maximum.

M. Dennefeld reports that a low-dispersion spectrum (range 370–715 nm, resolution 1 nm) was obtained with the 1.93-m telescope at the Haute Provence Observatory on Oct. 6.08 UT. The absence of hydrogen lines and the presence of a weak Si II 615-nm absorption line, together with various broad emissions, suggest a type Ia supernova ∼ 2 months past maximum.

Corrigendum. On IAUC 6088, the discovery date should have been given as Sept. 14 and the declination offset as 1′′.8 south.

1994 October 6 Brian G. Marsden